OpenStudy (anonymous):

A 560-gram salt-and-water solution contains 60 g of salt. This mixture is left in the open air, and 80 g of water evaporates from the solution.

OpenStudy (gorv):

welcome to OS :)

OpenStudy (gorv):

560 gram mixture salt =60 gram

OpenStudy (gorv):

water=560-60=500

OpenStudy (gorv):

evaporated=80 gram so water left= 500-80=420